# LLM-MedQA: Enhancing Medical Question Answering through Case Studies in Large Language Models

### v1

@@ Abstract
Accurate and efficient question-answering systems are essential for delivering high-quality patient care in the medical field. While Large Language Models (LLMs) have made remarkable strides across various domains, they continue to face significant challenges in medical question answering, particularly in understanding domain-specific terminologies and performing complex reasoning. These limitations undermine their effectiveness in critical medical applications. To address these issues, we propose a novel approach incorporating similar case generation within a multi-agent medical question-answering (MedQA) system. Specifically, we leverage the Llama3.1:70B model, a state-of-the-art LLM, in a multi-agent architecture to enhance performance on the MedQA dataset using zero-shot learning. Our method capitalizes on the model’s inherent medical knowledge and reasoning capabilities, eliminating the need for additional training data. Experimental results show substantial performance gains over existing benchmark models, with improvements of 7% in both accuracy and F1-score across various medical QA tasks. Furthermore, we examine the model’s interpretability and reliability in addressing complex medical queries. This research not only offers a robust solution for medical question answering but also establishes a foundation for broader applications of LLMs in the medical domain.
@@ I Introduction
Figure 1: Illustration of our proposed multi-agent architecture diagram, given a medical problem as an input to the larger model, which is divided into six phases: (1) Agent Generation, (2) Case Generation, (3) Analysis Proposition, (4) Report Generation, (5) Voting on the Decision, and (6) Making the Decision.
@@
In this study, we introduce a multi-agent framework, as shown in Fig. 1, to tackle the QA task on the medical domain. This framework leverages the Llama3.1:70B model, the off-the-shelf LLM model with 70 billion parameters, to enhance the performance of medical QA systems on the MedQA dataset [5]. Our multi-agent approach incorporates specialized agents to handle the inherent complexity of medical QA [6]. Each query in the system is assigned a series of experts, including question-specific analysis, option analysis, and case generation. A key innovation of our multi-agent system is the integration of a case generation module. This module autonomously generates supportive clinical cases tailored to the given question and selected options, as described in detail. Its purpose is to produce plausible and contextually accurate cases that substantiate the correct option for a specific medical problem. These cases, which are integrated into the reporting module, address a critical gap in existing systems by providing transparent and contextually rich explanations.
@@ III Methodology
In this section, we propose a specific multi-agent model framework to tackle the task of Medical Question Answering. The overall model consists of six components: (1) Multi-agent generation, which includes the creation of question experts and option experts. (2) Proposition analysis, this step involves a detailed analysis of the problem and the available options. (3) Case generation, which generates relevant cases based on the input questions and provided options. (4) Report digest, a report is generated by synthesizing insights from problem analysis, option analysis, and case generation. (5) Voting mechanism, where experts vote on the generated report, revising it as necessary if disagreements arise. This process continues iteratively until consensus is reached; and (6) Decision making, where the final report is used as the basis for selecting the correct answer. Apart from above components we described, we finally provide an overview of the algorithm used to facilitate expert voting and decision-making, and model selection explains why we chose to use the LLama 3.1:70B model.
@@ III-A Multi-Agent Generation
In the context of clinical medical problems, given a problem $q$ and a set of options $\text{op}=\{o_{1},o_{2},\ldots,o_{k}\}$, where $k$ denotes the total number of available options, the goal of this process is to assemble a team of experts. These include question experts, specialized in clinical problem analysis, $\text{questionExperts}=\{\text{qe}_{1},\text{qe}_{2},\dots,\text{qe}_{m}\}$ , as well as option experts specialized in analyzing the options, $\text{optionExperts}=\{\text{oe}_{1},\text{oe}_{2},\ldots,\text{oe}_{n}\}$, where $m$ and $n$ represent the respective numbers of question and option domain experts. Specifically, we assign a role to the model and provide instructions to guide it in generating the corresponding domain experts based on the input problems and options:
@@ III-B Proposition Analysis
Question Analyses: After consulting with the experts from relevant fields regarding the problem, we asked them to provide their individual analyses, which are then used to inform further reasoning. For each question $q$ and corresponding question expert $qe_{i}\in questionExperts$, we employ a large language model (LLM)to act as a domain-specific expert. Guided by the prompt $prompt_{qa}$, the LLM generates an analysis, represented by the following equation:
@@ IV-B Experiments Settings
In addition, we have chosen the Llama3.1:70B model due to its open-source nature, relatively low computational requirements, and cost-effectiveness compared to other models of similar performance. This choice ensures that the model remains accessible for research and application while providing high performance, making it particularly suitable for real-world healthcare QA scenarios. Additionally, the scalability and practicality of this model enhance the deployment of our approach in resource-constrained environments, ensuring both efficiency and effectiveness.
@@ IV-D Result and Analysis
Tab. I shows the overall results of comparison. It can be observed that our method achieves the best across all metrics, indicating that our multi-agent architecture offers advantages in real-world medical scenarios. In the few-shot setting, the Direct Inference method achieves an accuracy of 0.717, with Macro Precision, Recall, and F1-Score all at 0.717. Incorporating CoT results in a slight performance decrease, with accuracy dropping to 0.710 and macro metrics around 0.708–0.709. Adding SC to CoT further enhances performance, achieving the best results in the few-shot setting with an accuracy of 0.727 and macro metrics around 0.724. In the zero-shot setting, the baseline method performs relatively well with an accuracy of 0.714 and macro metrics of 0.715, while the CoT method alone demonstrates significant performance degradation, achieving only 0.698 in accuracy and 0.697 in macro metrics. The addition of SC to CoT shows notable improvement, achieving an accuracy of 0.719 and macro metrics of 0.718. The proposed ”Ours” method outperforms all other approaches in both zero-shot and few-shot settings, achieving the highest scores across all metrics, with an accuracy of 0.772 and macro metrics of 0.771. These results highlight the robustness and effectiveness of the proposed method, particularly in the challenging zero-shot scenario. The multi-agent architecture plays a key role in addressing complex decision-making problems. It not only integrates different expert opinions more effectively but also handles uncertainties common in healthcare scenarios, thereby improving the comprehensiveness and reliability of decision-making.

### v2

@@ Abstract
Accurate and efficient question-answering systems are essential for delivering high-quality patient care in the medical field. While Large Language Models (LLMs) have made remarkable strides across various domains, they continue to face significant challenges in medical question answering, particularly in understanding domain-specific terminologies and performing complex reasoning. These limitations undermine their effectiveness in critical medical applications. To address these issues, we propose a novel approach incorporating similar case generation within a multi-agent medical question-answering (MedQA) system. Specifically, we leverage the Llama3.1:70B model, a state-of-the-art LLM, in a multi-agent architecture to enhance performance of classification on the MedQA dataset using zero-shot learning. Our method capitalizes on the model’s inherent medical knowledge and reasoning capabilities, eliminating the need for additional training data. Experimental results show substantial gains over existing benchmark models, with improvements of 7% in both accuracy and F1-score across various medical QA tasks. Furthermore, we examine the model’s interpretability and reliability in addressing complex medical queries. This research not only offers a robust solution for medical question answering but also establishes a foundation for broader applications of LLMs in the medical domain.
@@ I Introduction
Figure 1: Illustration of our proposed multi-agent architecture diagram, given a medical problem as an input to the larger model, which is divided into six phases: (1) Agent Generation, (2) Case Generation, (3) Proposition Analysis, (4) Report Digest, (5) Voting Mechanism, and (6) Decision Making.
@@
In this study, we introduce a multi-agent framework, as shown in Fig. 1, to tackle the QA task on the medical domain. This framework leverages the Llama3.1:70B model, the off-the-shelf LLM model with 70 billion parameters, to enhance the performance of medical QA systems on the MedQA dataset [5]. Our multi-agent approach incorporates specialized agents to handle the inherent complexity of medical QA [6]. Each query in the system is assigned a series of experts, including question-specific analysis, option analysis, and case generation. A key innovation of our multi-agent system is the integration of a case generation module. This module autonomously generates supportive clinical cases tailored to the given question and selected options. Its purpose is to produce plausible and contextually accurate cases that substantiate the correct option for a specific medical problem. These cases, which are integrated into the reporting module, address a critical gap in existing systems by providing transparent and contextually rich explanations.
@@ III Methodology
In this section, we propose a specific multi-agent model framework (Fig. 1) to tackle the task of Medical Question Answering. The overall model consists of six components: (1) Multi-agent generation, which includes the creation of question experts and option experts. (2) Proposition analysis, this step involves a detailed analysis of the problem and the available options. (3) Case generation, which generates relevant cases based on the input questions and provided options. (4) Report digest, a report is generated by synthesizing insights from problem analysis, option analysis, and case generation. (5) Voting mechanism, where experts vote on the generated report, revising it as necessary if disagreements arise. This process continues iteratively until consensus is reached; and (6) Decision making, where the final report is used as the basis for selecting the correct answer. Apart from above components we described, we finally provide an overview of the algorithm used to facilitate expert voting and decision-making, and model selection explains why we chose to use the LLama 3.1:70B model.
@@ III-A Multi-Agent Generation
In the context of clinical medical problems, given a problem $q$ and a set of options $\text{op}=\{o_{1},o_{2},\ldots,o_{k}\}$, where $k$ denotes the total number of available options, the goal of this process is to assemble a team of experts. These include question experts, specialized in clinical problem analysis, $\text{questionExperts}=\{\text{qe}_{1},\text{qe}_{2},\dots,\text{qe}_{m}\}$ , as well as option experts specialized in analyzing the options, $\text{optionExperts}=\{\text{oe}_{1},\text{oe}_{2},\ldots,\text{oe}_{n}\}$, where $m$ and $n$ represent the respective numbers of question and option domain experts. Specifically, we assign a prompt to the model and provide instructions to guide it in generating the corresponding domain experts based on the input problems and options:
@@ III-B Proposition Analysis
Question Analyses: After consulting with the experts from relevant fields regarding the problem, we asked them to provide their individual analyses, which are then used to inform further reasoning. For each question $q$ and corresponding question expert $qe_{i}\in questionExperts$, we employ a LLM to act as a domain-specific expert. Guided by the prompt $prompt_{qa}$, the LLM generates an analysis, represented by the following equation:
@@ IV-B Experiments Settings
In addition, we have chosen the Llama3.1:70B model due to its open-source nature, relatively low computational requirements, and cost-effectiveness compared to other models of similar performance. This choice ensures that the model remains accessible for research and application while providing high performance, making it particularly suitable for real-world healthcare QA scenarios.
@@ IV-D Result and Analysis
Tab. I shows the overall results of comparison. It can be observed that our method achieves the best across all metrics, indicating that our multi-agent architecture offers advantages in real-world medical scenarios. In the few-shot setting, the Direct Inference method achieves an accuracy of 0.717, with Macro Precision, Recall, and F1-Score all at 0.71. Incorporating CoT results in a slight performance decrease, with accuracy dropping to 0.710 and macro metrics around 0.708. Adding SC to CoT further enhances performance, achieving the best results in the few-shot setting with an accuracy of 0.727 and macro metrics around 0.724. In the zero-shot setting, the baseline method performs relatively well with an accuracy of 0.714 and macro metrics of 0.715, while the CoT method alone demonstrates significant performance degradation, achieving only 0.698 in accuracy and 0.697 in macro metrics. The addition of SC to CoT shows notable improvement, achieving an accuracy of 0.719 and macro metrics of 0.718. The proposed ”Ours” method outperforms all other approaches in both zero-shot and few-shot settings, achieving the highest scores across all metrics, with an accuracy of 0.772 and macro metrics of 0.771. These results highlight the robustness and effectiveness of the proposed method, particularly in the challenging zero-shot scenario. The multi-agent architecture plays a key role in addressing complex decision-making problems. It not only integrates different expert opinions more effectively but also handles uncertainties common in healthcare scenarios, thereby improving the comprehensiveness and reliability of decision-making.